# Reconstructing Hand-Held Objects in 3D

### v2

@@ 6 Experiments
Section 6.1 details the datasets that were used for training and evaluation. Section 6.2 defines the evaluation metrics used for comparisons. Section 6.3 quantifies the advantages of using MCC-HO, e.g., a hand-centric approach, in comparison to the original MCC model [85]. We also quantitatively (Section 6.4) and qualitatively (Section 6.5) evaluate our model against existing work in model-free, monocular hand-object reconstruction [24, 90] and find that MCC-HO achieves state-of-the-art performance on all three datasets used [5, 3, 38]. In particular, the methods most similar to ours are HO [24] and IHOI [90]. The quality of the MCC-HO predicted geometry is also shown to be sufficient for accurate rigid alignment of template objects using ICP. Also in Section 6.5.1, we provide qualitative results for our Retrieval-Augmented Reconstruction (RAR) approach on in-the-wild Internet images to demonstrate the feasibility and scalability of our automated model retrieval and fitting. Specifically, we apply MCC-HO and RAR to unlabeled images from the 100DOH dataset [66].
6.1 Datasets
HOI4D: The HOI4D dataset [38] consists of 4,000 lab videos of hands interacting with specified categories of common objects. We selected all sequences containing one of 6 hand-held object categories, and divide the sequences into training, validation and test sets. In total, we used 1142 training sequences, 6 validation sequences, and 12 test sequences consistent with DiffHOI [91]. For the validation set, we selected one video per category and favored object instances that appear the least frequently in the dataset.
DexYCB: The DexYCB dataset [5] contains 1,000 lab videos of hands interacting with 21 YCB objects. We used the “s0” dataset split, which contains 800 training, 40 validation, and 160 test sequences (consistent with prior work [7, 6]).
MOW/RHOV: MOW [3] is a dataset of 512 images from 100 Days of Hands [66] labeled with 3D hands and objects. We use the same splits as [90] with 350 training images and 92 test images. Additionally, we use 110 videos from RHOV [51], each of which is a 2 second clip centered at the corresponding MOW frame. In the RHOV dataset, 19 videos overlap with the MOW test set and thus we do not use them during training.
100DOH: The 100 Days of Hands dataset [66] contains 131 days of footage collected from Internet videos showing hands interacting with objects. The vast majority of videos do not have associated 3D hand or object labels, so we evaluate how well our approach generalizes using examples from this unlabeled subset.
6.2 Evaluation Metrics
Since the focus of our approach is on the object manipulated by the hand, we quantitatively evaluate object reconstruction quality using standard 3D metrics, Chamfer Distance (CD) and F-score reported at 5mm and 10mm. Both metrics are computed on point sets, and thus the point clouds inferred by MCC-HO are directly used. The 3D hand provided by each dataset is used as input; note that MOW uses FrankMocap [63] and HOI4D optimizes hand pose using 2D keypoint annotations. For each example, we sample 10k points on the surface of the ground truth mesh consistent with [90, 91] (no transformations are applied to the ground truth geometry). The MCC-HO predicted point clouds are supersampled/subsampled to obtain 10k points as well. In order to account for unknown ground truth camera parameters, ICP with scaling [58] is used to align all predicted geometry with the ground truth mesh. Median metrics are reported for each test dataset.
Chamfer Distance: The Chamfer Distance metric we use between two point sets $A,B\in\mathbb{R}^{N\times 3}$ is:
[…]
F-score: For 3D evaluation of predicted points, the F-score at a given threshold $t$ is defined as:
$F_{1}(X_{pd},X_{gt})=2\vspace{1mm}\frac{\text{Precision}(X_{pd},X_{gt},t)\cdot\text{Recall}(X_{pd},X_{gt},t)}{\text{Precision}(X_{pd},X_{gt},t)+\text{Recall}(X_{pd},X_{gt},t)}.$ (7)
6.3 MCC Baseline
Because our model is built on top of MCC [85], we begin by evaluating whether MCC-HO leads to performance gains for the task of hand-held object reconstruction. For MCC, we use the released model that was trained on the CO3Dv2 dataset [60] and apply an existing monocular depth estimator [57] to obtain depth maps (as MCC takes an RGB-D image as input). Table 1 compares 3D metrics between MCC and MCC-HO computed on the MOW test set. We find that fine-tuning MCC for the specific task of hand-held object reconstruction improves performance on test images unseen by both models.
Model Parameters F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
MCC [85] 0.10 0.21 31.6
MCC-HO 0.15 0.31 15.2
[…]
@@ 6.3 MCC Baseline
MCC-HO 0.36 0.60 3.74 0.15 0.31 15.2 0.52 0.78 1.36
[…]
@@ 6.4 Quantitative Evaluation
MCC-HO is quantitatively evaluated on a variety of labeled hand-object datasets and in comparison to existing model-free reconstruction approaches [24, 90]. See Table 2. In order to sample points more densely in hand-object regions, the dimensions and granularity of the voxel grid used to query MCC-HO are determined from an initial coarse prediction using a default voxel grid. For DexYCB, we use every 5th frame of the “s0” test split, and for HOI4D, we use all frames from the 12 test sequences that contain a ground truth 3D hand label. The results for IHOI [90] and HO [24] evaluated on the MOW dataset are obtained from the IHOI and DiffHOI [91] papers. Because AlignSDF [7] and gSDF [6] trained DexYCB-specific models, whereas we train a single model on multiple datasets, we do not compare to their published metrics.
Our method achieves state-of-the-art performance for the Chamfer Distance and F-score metrics on all three datasets, implying that our technique is not overfitted to any particular dataset. We conjecture that increasing the number of unique object instances seen during training is important to model generalization. In addition, the pretrained MCC model provides a strong object prior for initialization.
@@ 6.5 Qualitative Evaluation
We qualitatively evaluate (1) the reconstruction quality of MCC-HO and (2) how well template objects can be rigidly aligned to the predicted point clouds in 3D (a necessary and sufficient condition for downstream tasks). Figure 5 shows predicted point clouds for a number of MOW test examples, as well as the result of rigidly aligning the ground truth CAD model with each point cloud using ICP. Since MCC-HO is trained with monocular images, the predicted RGB colors are only valid when viewed from the given camera. Because the object geometry is conditioned on the 3D hand, our approach tends to predict reasonable hand-object contact. Additionally, the object in contact with the hand is correctly identified by MCC-HO despite being provided a hand-object bounding box and no segmentation maps.
Figure 5: MCC-HO results for MOW test examples. The input image (top, left), network-inferred hand-object point cloud (top, right), rigid alignment of the ground truth mesh with the point cloud using ICP (bottom, left), and an alternative view of the point cloud (bottom, right) are shown.
Figure 6: Our approach applied to test images from MOW (Rows 1-3), DexYCB (Row 4), and HOI4D (Row 5). Each example includes the input image (left), rigid alignment of the ground truth mesh with the MCC-HO predicted point cloud using ICP (middle), and an alternative view (right).
@@
Figure 6 shows additional visualizations of applying rigid alignment of the template model to MCC-HO results. We include held-out test examples from all three datasets used for training: MOW, DexYCB, and HOI4D. Because MCC-HO is fine-tuned from a general object reconstruction transformer model (MCC), it is robust to unseen object instances at test time. For instance, objects in the MOW test images were never seen by MCC-HO during training. From a practical standpoint, we find 3D-to-3D fitting of template objects with the MCC-HO predicted geometry (e.g., via ICP) to be significantly more efficient than prior 3D-to-2D or per-example optimization approaches [3, 51, 91]. Moreover, optimizing 3D pose with respect to 2D images tends to get stuck in local minima. As Figure 6 demonstrates, this 3D fitting results in physically plausible hand-object interactions that when projected align with the image data.
6.5.1 Retrieval-Augmented Reconstruction.
The ultimate goal of RAR is to provide a scalable paradigm for creating large hand-object 3D datasets; to this end, we explore the potential to scale up our approach on unlabeled images. Specifically, we take 9 images from 100DOH [66] that do not have 3D ground truth (e.g., not in MOW) and apply MCC-HO combined with RAR to reconstruct hand-object 3D geometry. See Figure 7. Estimated 3D hands for each image are obtained using HaMeR [52]. At test time, our method does not require an input hand-object segmentation mask [3, 51], depth image [85], or hand joint locations [90, 8]. Rather, we simply run feed-forward inference using MCC-HO, retrieve a 3D object mesh using RAR, and run ICP to align this mesh with the inferred point cloud. As more training data is collected through RAR and human-in-the-loop verification, we expect MCC-HO to naturally improve as well (similar in spirit to the data collection process used in Segment Anything [29]).
6.6 Limitations
While our 3D-to-3D rigid alignment approach using ICP is efficient, some misalignment may occur when rendering the reconstruction from the input view camera. This is because we do not use image-based cues for alignment, but note that this can be easily fixed via inverse rendering as in [3, 51]. In particular, the 6DOF object pose can be fine-tuned by minimizing the silhouette loss and hand-object contact/collision losses defined in [3]. See Figure 8. Regardless, our reconstruction provides a more automatic/robust initialization of the 3D object model than manual or random initialization, as was done in prior work [3, 51]. In addition, because MCC-HO is conditioned on an input 3D hand, the accuracy of the predicted hand pose naturally affects object reconstruction performance.
Figure 8: The input image is not used when running ICP, which can sometimes result in image misalignment errors (Row 1). A postprocess can be applied to resolve silhouette misalignment and any hand-object contact issues (Row 2).
@@ 7 Conclusion
In this paper, we present a novel paradigm for reconstructing hand-held objects that combines the respective benefits of model-free and model-based prediction. Our transformer-based model, MCC-HO, is trained to predict hand-object geometry given a single RGB image and estimated 3D hand. Experiments conducted using the DexYCB, MOW, and HOI4D datasets all demonstrate that MCC-HO achieves state-of-the-art performance in hand-held object reconstruction. We additionally present Retrieval-Augmented Reconstruction (RAR), an automatic method for object model retrieval that leverages recent advances in large language/vision models and 3D object datasets. These two approaches can be combined to scale the amount of labeled hand-object interaction data, as suggested by our results using unlabeled 100DOH images [66]. The combination of MCC-HO and RAR can enable the construction of large hand-object datasets at-scale; for future work, we plan to expand on this paradigm to label a larger dataset of Internet videos.

### v3

@@ 7 Experiments
Section 7.1 details the datasets that were used for training and evaluation. Section 7.2 defines the evaluation metrics used for comparisons. We also quantitatively (Section 7.3) and qualitatively (Section 7.4) evaluate against existing work and find that our approach achieves state-of-the-art performance on all three datasets used [4, 2, 41]. Also in Section 7.4.1, we provide qualitative results for our Retrieval-Augmented Reconstruction (RAR) approach applied to in-the-wild Internet videos from 100DOH [71] to demonstrate the feasibility and scalability of our automated model retrieval and fitting. Section 7.5 includes ablation studies for MCC-HO, RAR, and each component of our Viterbi temporal alignment algorithm.
@@
HORSE [58] 0.20 0.41 7.16 0.08 0.20 27.2 0.21 0.41 7.68
MCC-HO (Ours) 0.36 0.60 3.74 0.15 0.31 15.2 0.52 0.78 1.36
Table 1: We compare our method, MCC-HO, to prior works on held-out test images from DexYCB, MOW, and HOI4D. Chamfer Distance (cm2) and F-score (5mm, 10mm) are reported.
Method F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
DiffHOI [89] 0.62 0.91 0.8
G-HOP [96] 0.76 0.97 0.4
MCC-HO + RAR (Ours) 0.91 0.98 0.28
Table 2: Quantitative evaluation of video-based hand-held object reconstruction methods on the HOI4D test dataset. Mean metrics are reported for consistency with prior work.
7.1 Datasets
HOI4D. The HOI4D dataset [41] consists of 4,000 lab videos of hands interacting with specified categories of common objects. We selected all sequences containing one of 6 hand-held object categories, and divide the sequences into training, validation and test sets. In total, we used 1142 training sequences, 6 validation sequences, and 12 test sequences consistent with DiffHOI [95]. For the validation set, we selected one video per category and favored object instances that appear the least frequently in the dataset.
DexYCB. The DexYCB dataset [4] contains 1,000 lab videos of hands interacting with 21 YCB objects. We used the “s0” dataset split, which contains 800 training, 40 validation, and 160 test sequences.
MOW/RHOV. MOW [2] is a dataset of 512 images from 100 Days of Hands [71] labeled with 3D hands and objects. We use the same splits as [94] with 350 training images and 92 test images. Additionally, we use 110 videos from RHOV [55], each of which is a 2 second clip centered at the corresponding MOW frame. In the RHOV dataset, 19 videos overlap with the MOW test set and thus we do not use them during training.
100DOH. The 100 Days of Hands dataset [71] contains 131 days of footage collected from Internet videos showing hands interacting with objects. The vast majority of videos do not have associated 3D hand or object labels, so we evaluate how well our approach generalizes using examples from this unlabeled subset.
7.2 Evaluation Metrics
Since the focus of our approach is on the object manipulated by the hand, we quantitatively evaluate object reconstruction quality using standard 3D metrics: CD (cm2) and F-score reported at 5mm and 10mm. Both metrics are computed on point sets. The 3D hand provided by each dataset is used as input; note that MOW uses FrankMocap [69] and HOI4D optimizes hand pose using 2D keypoint annotations. For each example, we sample 10k points on the surface of the ground truth mesh consistent with [94, 95] (no transformations are applied to the ground truth geometry). The MCC-HO predicted point clouds are supersampled/subsampled to obtain 10k points as well. In order to account for unknown ground truth camera parameters, ICP with scaling [64] is used to align all predicted geometry with the ground truth mesh. Median metrics are reported unless otherwise noted.
Chamfer Distance. The CD metric we use between two point sets $A,B\in\mathbb{R}^{N\times 3}$ is:
where the point sets should have the same size. This equation is consistent with prior work [5, 6, 94, 95]. We report CD in cm2 as in[5, 6].
F-score. The F-score at a given threshold $t$ is defined as a function of precision $P$ and recall $R$:
$F_{1}(X_{pd},X_{gt})=2\vspace{1mm}\frac{P(X_{pd},X_{gt},t)\cdot R(X_{pd},X_{gt},t)}{P(X_{pd},X_{gt},t)+R(X_{pd},X_{gt},t)}.$ (9)
@@ 7.3 Quantitative Evaluation
We quantitatively evaluate MCC-HO on a variety of labeled hand-object datasets and in comparison to existing model-free reconstruction approaches [23, 94, 58]. See Table 1. In order to sample points more densely in hand-object regions, the dimensions and granularity of the voxel grid used to query MCC-HO are determined from an initial coarse prediction using a default voxel grid. For DexYCB and HOI4D, we use every 5th frame of each test video sequence. The results for IHOI [94] and HO [23] evaluated on the MOW dataset are obtained from the IHOI and DiffHOI [95] papers; otherwise, we used available public repositories of prior work. Because AlignSDF [5] and gSDF [6] trained DexYCB-specific models, whereas we train a single model on multiple datasets, we do not compare to their published metrics. Our method achieves state-of-the-art performance for the Chamfer Distance and F-score metrics on all three datasets, implying that our technique is not overfitted to any particular dataset. We conjecture that increasing the number of unique object instances seen during training is important to model generalization. In addition, the pretrained MCC model provides a strong object prior for initialization.
We also compare to existing video-based approaches for RGB hand-held object reconstruction in Table 2, noting that the ground truth object is used for our RAR rigid alignment (a reasonable comparison as DiffHOI [95] and GHOP [96] train category-specific diffusion models using the ground truth HOI4D 3D objects).
@@ 7.4 Qualitative Evaluation
We qualitatively evaluate (1) the reconstruction quality of MCC-HO and (2) how well retrieved object models using RAR can be rigidly aligned to the predicted point clouds in 3D (a necessary and sufficient condition for downstream tasks). Figure 4 shows predicted point clouds for a number of MOW test examples, as well as the result of rigidly aligning the ground truth CAD model with each point cloud. Since MCC-HO is trained with monocular images, the predicted RGB colors are only valid when viewed from the given camera. Because the object geometry is conditioned on the 3D hand, our approach tends to predict reasonable hand-object contact. Additionally, the object in contact with the hand is correctly identified by MCC-HO without input segmentation maps.
Figure 4: MCC-HO results for MOW test examples. The input image (top, left), network-inferred hand-object point cloud (top, right), rigid alignment of the ground truth mesh with the point cloud (bottom, left), and an alternative view of the point cloud (bottom, right) are shown.
@@
7.4.1 Retrieval-Augmented Reconstruction
The ultimate goal of RAR is to provide a scalable paradigm for creating large hand-object 3D datasets; to this end, we explore the potential to scale up our approach on unlabeled images and videos. Figure 5 shows our results for 9 images from 100DOH [71] that do not have 3D ground truth (e.g., not in MOW). Estimated 3D hands for each image are obtained using HaMeR [56]. At test time, our method does not require an input hand-object segmentation mask [2, 55], depth image [89], or hand joint locations [94, 8]. Please see the supplementary materials for additional results.
7.5 Ablation Studies
MCC Baseline. Because our model is built on top of MCC [89], we begin by evaluating whether MCC-HO leads to performance gains for the task of hand-held object reconstruction. For MCC, we use the released model that was trained on the CO3Dv2 dataset [67] and apply an existing monocular depth estimator [63] to obtain depth maps (as MCC takes an RGB-D image as input). Table 3 compares 3D metrics between MCC and MCC-HO computed on the MOW test set. We find that using MCC-HO, which was trained on the specific task of hand-held object reconstruction, leads to performance improvements on test images.
Method F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
MCC [89] 0.10 0.21 31.6
MCC-HO 0.15 0.31 15.2
Table 3: Analysis of how hand-held object reconstruction benefits from conditioning on the hand using the MOW test dataset.
RAR Success Rate. In order to quantify the success rate of RAR (e.g., whether a generated 3D object matches the input image), we use GPT-4(V) and Genie to automatically generate objects for all 512 images in the MOW dataset. Genie is able to generate an object 82% of the time, and otherwise returned an error. The results are included in the supplementary materials.
Video Alignment Components. We aim to understand the contributions of each cost term in our Viterbi algorithm, e.g., MCC-HO and DINOv2. See Table 4. In order to best assess the quality of rigid alignment, the ICP postprocess described in Section 7.2 is not applied. The results demonstrate that both terms contribute to the efficacy of our Viterbi algorithm; in practice, we observe that using both 2D and 3D cues is particularly important for objects that lie outside the training distribution of MCC-HO (DexYCB consists of commonplace YCB objects).
Viterbi Cost Terms F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
MCC-HO 0.26 0.52 4.18
DINOv2 0.23 0.45 6.38
MCC-HO + DINOv2 0.27 0.52 4.10
Table 4: Analysis of how each of our Viterbi cost terms contributes to the final temporal alignment on the DexYCB test dataset.
@@ 8 Conclusion
In this paper, we present a novel paradigm for reconstructing hand-held objects that combines the respective benefits of model-free and model-based prediction. Our transformer-based model, MCC-HO, is trained to predict hand-object geometry given a single RGB image and estimated 3D hand. Experiments conducted using the DexYCB, MOW, and HOI4D datasets all demonstrate that MCC-HO achieves state-of-the-art performance in hand-held object reconstruction. We additionally present Retrieval-Augmented Reconstruction (RAR), an automatic method for object model retrieval that leverages recent advances in large language/vision models and 3D object datasets. These two approaches can be combined to scale the amount of labeled hand-object interaction data, as suggested by our results using unlabeled 100DOH videos [71].
Limitations. As 3D generative models improve, RAR will naturally be able to retrieve more accurate and realistic geometry. Our approach also relies on the ability to obtain accurate hand and object masks which are tracked across frames. We expect that both MCC-HO and our rigid alignment algorithm will reap the benefits of any future image segmentation/tracking improvements.